If the velocity of money (V) and real output (Q) were increasing at approximately the same rate, then:


A. it would be impossible for monetary authorities to control inflation.

B. monetary acceleration would not lead to inflation.

C. inflation would be closely related to the long-run rate of monetary expansion.

D. the money supply would not need to increase in order to maintain stable prices as output rose over time

E. both c. and d. would be true.
Answer #1

Solution:

Quantity theory of money states that: M*V = P*Q; M is the money supply, V is the velocity of money, P is the price and Q is the real output.

Then, M/P = Q/V

In terms of growth, then, growth rate of M - growth rate of P = growth rate of Q - growth rate of V

With growth rate of Q approximately equal to growth rate of V, we have

Growth rate of M - Growth rate of P = 0

Growth rate of M = Growth rate of P

Also, we know that growth rate in price level is the inflation rate, and by monetary expansion we mean that growth rate of money supply be positive. Then, using the expression above we can conclude two points:

1) Inflation and monetary expansion growth rate are directly proportional, hence they indeed are closely related.

2) To keep inflation rate 0, that is change in price level 0 (which is maintaining stable prices), change in money supply shall also be 0 (that is money supply would not be increased)

Hence, the correct option is E) both (C) and (D) would be true.